A charge of 2.15 mC is placed at each corner of a square 0.240 m on a side.

Part A

Determine the magnitude of the force on each charge.

Express your answer to three significant figures and include the appropriate units.

PLEASE DRAW OUT AND FULLY EXPLAIN YOUR STEPS AND ANSWERS. i keep getting my answers wrong and I cant figure out where I am messing up. I WILL RATE AND COMMENT!!!!!

THANKS
Answer #1

force due to adjacent charge

F1 = F3 = k q1 q / a^2 = 9*10^9* (2.15*10^-3)^2 / 0.24^2

F1 = F3 = 7.22*10^5 N

resultant force of F1andF3

F13^2 = F1^2 + F3^2

F13 = 10.21*10^5 N

force due to corner charge

F2 = k q^2 / ( 2 a^2)

F2= 9*10^9*( 2.15*10^-3)^2 / ( 2* 0.24^2)

F2 = 3.61*10^5 N

net force on any charge,

F = F13 + F2

F = ( 10.21 + 3.61) *10^5

F = 13.82*10^5 N

========

Comment in case any doubt, will reply for sure. Goodluck